OpenStudy (anonymous):

Divide 35b5 + 20ab3 + 20a2b2 by 5b2. What is the quotient?

OpenStudy (anonymous):

7b^3 + 4ab + 4a^2

OpenStudy (ranga):

\[\Large \frac{ 35b ^{5} + 20ab^{3} + 20a^{2}b^{2} }{ 5b^{2} } = \frac{ 35b^{5} }{ 5b^{2} } + \frac{ 20ab^{3} }{ 5b^{2} } + \frac{ 20a^{2}b^{2} }{ 5b^{2} }\]